OpenStudy (anonymous):

calculate the density of a rectangular solid which has a mass of 25.71g. it is 2.30cm long, 4.01 cm wide, and 1.82 cm high.

OpenStudy (anonymous):

find the volume of the solid = 2.3 * 4.01 * 1.82 cm^3 density = mass/volume mass = 25.71g put this in the formula and get your answer in unit g/cm^3

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Mass=25.7gm dimensions, l=2.3cm b=4.01cm h=1.82cm Thus,Density=Mass/Volume=m/lbh=25.7/2.3*4.01*1.82 g cm-3